# A stochastic reconstruction theorem on rectangular increments with an application to a mixed hyperbolic SPDE

### v2

@@ Abstract
We extend the stochastic reconstruction theorem to a setting where the underlying family of distributions satisfies some natural conditions involving rectangular increments. This allows us to prove the well-posedness of a new class of mixed stochastic partial differential equations of hyperbolic type which combines standard Walsh stochastic integration and Young products.
@@ 1.1 Stochastic reconstruction and hyperbolic SPDEs
where $\lambda\in(0,1]$, $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, $\psi_{\mathbf{x}}^{\lambda}(\cdot)=\lambda^{-d}\psi(\lambda^{-1}(\cdot-\mathbf{x}))$ is a rescaled and recentered version of a test function $\psi$ and $\lesssim$ denotes inequality up to a constant. Further analytic conditions were studied in [BL23, ZK22].
In this paper, we are interested in a generalization of this result where the underlying family $F$ is composed of random distributions depending on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ which are properly adapted to a multiparameter filtration =$\mathcal{F}=\{\mathcal{F}_{\mathbf{x}}\colon\mathbf{x}\in\mathbb{R}^{d}\}$ (see Definition 2.15). This setting is not novel in the literature. A previous result [Ker23] showed we can modify the conditions (1.1) into
@@
where $\delta>0$, $\|X\|_{m}=\mathbb{E}(|X|^{m})^{1/m}$ and $\mathbb{E}^{i}_{\mathbf{x}}(X)=\mathbb{E}[X|\mathcal{F}^{i}_{\mathbf{x}}]$ is the conditional expectation with respect to $\mathcal{F}^{i}_{\mathbf{x}}$, the filtration generated by the $i$-th component of $\mathbf{x}$. Remarkably, the uniqueness result holds even in the case $20>\gamma>-d/$ 2 provided that $\gamma+\delta>0$. Heuristically, we leverage the gain of regularity obtained by conditioning the germ by allowing a lower index of general regularity $\gamma$. This type of argument is not new in the rough analysis literature but it follows the same idea of the stochastic sewing lemma [Le20], from which [Ker23] was inspired. Moreover, stochastic sewing arises as a special case of the stochastic reconstruction when $d=1$, see [Ker23, Section 5]. From a stochastic analysis perspective, stochastic reconstruction provides a new analytic framework to study Walsh integration [Wal86] with respect to a white noise $\xi$ over +$\mathbb{R}^{d}_{+}:=[0,+\infty)^{d}$ as a Young product, see [CZ20]. Recalling the existence of a measurable modification of $\xi$ in the space of distribution, for any adapted process $Y$ satisfying some weak regularity assumption one can check that the germ $(Y\cdot\xi)_{\mathbf{x}}(\psi):=Y_{\mathbf{x}}\xi(\psi)$ satisfies (1.2) and one obtains
@@
where the left-hand side is a standard Walsh integral, see [Ker23, Theorem 53].
@@
where ${\bm{\lambda}}=(\lambda_{1}\,,\ldots\,,\lambda_{d})\in(0,1]^{d}$, $\mathbf{x}=(x_{1}\,,\ldots,x_{d})$ $\mathbf{y}=(y_{1}\,,\ldots,y_{d})$ are in compact set, and $\psi_{\mathbf{y}}^{{\bm{\lambda}}}$ is a rescaled test function like in (2.7). Then there exist unique family of distributions $\mathcal{R}^{\theta}_{\mathbf{x}}(F)$ for each $\theta\subset\{1,\dots,d\}$ , such that $\mathcal{R}_{\mathbf{x}}^{\theta}(F)$ does not depend on $x_{i}$ for $i\in\theta$ and for any non empty subset one has the growth $\eta\subset\theta\neq\emptyset$,
@@
Unlike the statements in [CZ20, Ker23], in this statement we allow one to freeze certain variables $x_{i},i\notin\theta\subset\{1,\dots,d\}$ and only apply the reconstruction techniques for directions $i\in\theta$. As a result, we do not get a single distribution $\mathcal{R}(F)$, but a family of partial reconstructions $\mathcal{R}^{\theta}_{\mathbf{x}}(F)$. Let us illustrate these on example (1.3) in $d=2$. By the same arguments as [Ker23, Theorem 53],
@@
It should be noted that the proof of the stochastic reconstruction theorem is closer to the original proof of Martin Hairer than the later improvements [CZ20, BL23, ZK22]. This is because Martin Hairer’s proof is based on wavelet approximations, which approximate the target distribution $f$ with the sum
@@
where $\phi^{\mathbf{n}}_{\mathbf{y}}$ is a Daubechies-wavelet and $\Delta_{\mathbf{n}}$ is an increasingly fine mesh over $\mathbb{R}^{d}$, see Section 2. This particular structure allows to treat the convergence of these sums in the same way as [Le20] by using a proper estimate on sums via a generalized BDG inequality, see Lemma 2.14. Moreover, wavelets also play an important role when we want to integrate the reconstruction to obtain a proper random field (see Proposition 4.7).
Coming back again to dimension one, the possibility of multiplying a Brownian motion with a deterministic distribution via stochastic sewing was one of the inspiring principles to develop in [FHL21] a complete theory of rough stochastic differential equations of the form
$dY_{t}=b_{t}(Y_{t})+\sigma_{t}(Y_{t})dW_{t}+f_{t}(Y_{t})d\mathbf{X}_{t}$ (1.7)
[…]
@@
where $\xi$ is a white noise over $\mathbb{R}_{+}^{d}$ and $Z\colon\mathbb{R}_{+}^{d}\to\mathbb{R}$ is a deterministic Hölder field in $C^{\beta}$, see Section 3.1. Moreover, $\sigma\colon\mathbb{R}\to\mathbb{R}$ is a generic non-linearity and $f\colon\Omega\times\mathbb{R}_{+}^{d}\to\mathbb{R}$ is a given random field. The reconstruction theorem in Theorem 1.1 allows us to provide a general solution theory for (1.8).
Formally speaking, equation (1.8) represents a simplified version of (1.7) in several variables but at the same time it is connected with new and old problems both in PDE and SPDE theory: firstly, when $\sigma=0$ there is no randomness, the equation (1.8) is a special case of a differential equation driven by a Hölder field on hyper-cubes, see [Har21] or a special rotated wave equation when $d=2$, see [QST07]. The choice of this specific non-linearity plays an important role in the theory of signature Kernels, see e.g. [SCF+21, LL24]. Concerning the opposite case when $f=0$ the underlying SPDE is classical and it has been intensively studied in dimension $2$, see e.g. [Cai72, NY89b, NY89a, MSS94]. To our knowledge, equations such as (1.8) do not seem to have been addressed in the literature yet, and from a technical point of view, both classical deterministic techniques and techniques based on standard stochastic integration fail to establish well-posedness of this problem.
To give a rigorous meaning to equation (1.8), we recall the identification of stochastic integration with the reconstruction in (1.3) and we will describe the products in the right-hand side as reconstruction of germs of the simple form recalled before. Then we can write the equation into its mild solution form after applying an integration map
@@
Recently, a new stochastic extension of the multiparameter sewing lemma was proposed in [BHK23] to study the effect of regularisation by noise phenomena on a similar SPDE. While some techniques developed in that article are fundamental in the proof of the reconstruction theorem, we think that direct use of that result to solve (1.8), even if possible, would require a higher complexity in the study of the spaces $C^{\alpha,\delta}L_{m}$. In particular, with stochastic reconstruction, one has the advantage of not dealing with multiparameter delta operators, which become cumbersome in high dimensions. In our context, we treated this complexity by developing a specific integration theory in Section 4. Furthermore, a formulation of this SPDE using a reconstruction theorem could be more helpful in extending techniques inspired by regularity structures to more general hyperbolic SPDEs. Further details between our version of stochastic reconstruction and the stochastic multiparameter sewing lemma are studied in Section 5.
@@ 1.2 Outline of the paper
In Section 5, we show that the multiparameter stochastic sewing lemma from [BHK23] is a special case of our reconstruction theorem 3.12. This is shown in Proposition 5.4. It is based on the same construction, which connects sewing and reconstruction (see e.g. [BZ22]), as well as stochastic sewing and stochastic reconstruction (see [Ker23]), in dimension $d=1$.
@@ 1.3 Open extensions arising from our work
We conclude the introduction by outlining two natural extensions raised by the results we established.
In our current formulation, we can only solve (1.9) for non-linearities of the form $f(\mathbf{x})u(\mathbf{x})$. One natural extension of the problem would be to consider a generic non-linearity $g(u)$ in (1.8). However, it seems that the structure of the spaces $C^{\alpha,\delta}L_{m}$ is not well suited to take generic compositions $g(u)$ for $u\in C^{\alpha,\delta}L_{m}$ into account. That is, even for smooth functions $g$, $g(u)$ might not belong to the same space as $u$. This stands in stark contrast to the deterministic case: In $d=2$, it is well-known that if $g\in C^{2}$ then $g(u)\in C^{\bm{\alpha}}$ for any $u\in C^{\bm{\alpha}}$ and ${\bm{\alpha}}=(\alpha_{1},\alpha_{2})\in(0,1)^{2}$ (see e.g. [QST07, Lemma 3.1]). Moreover, for any Hölder function $\sigma\in C^{\eta}$ for some $\eta\in(1,2)$, we still have that $\sigma(u)\in C^{\tilde{\alpha}}$ for a some $\tilde{\alpha}<\alpha$ (see [BHR22, Lemma 15]). Both of these results rely on the application of the classical Taylor formula to the rectangular increment
@@
In [FHL21], the authors encountered a similar problem when $d=1$ which they managed to solve with the use of $L_{m,n}$ norms. Using our notations, this would be equivalent to controlling the quantities
@@ Acknowledgements
CB is supported by the ERC starting grant “LoRDeT”; he also acknowledges the support of TU Berlin where he was employed when this project was started. HK is supported by the Research Unit FOR2402 in TU Berlin and the International Research Training Group 2544 of the Berlin Mathematical School. Both authors would like to thank the organizers of the conference “Structural Aspects of Signatures and Rough Paths” held at the Centre for Advanced Study of Oslo, where a preliminary version of these results was presented.
@@
$(\mathbf{y}^{\mathbf{x}})_{\theta}=\prod_{i\in\theta}y_{i}^{x_{i}}\,,$
@@ 2.2 Wavelet bases
To prove the reconstruction theorem in the rectangular increment setting we will have to use all the regularising effects of wavelet theory, going back to [Dau88]. First, we will recall some general facts about wavelet basis, mainly following [Hai14, Ker23, Mey93].
@@
where we define $\hat{\phi}^{\zeta,\mathbf{n}}_{\mathbf{x}}$ analogously. Using the notation $\Delta_{n}:=\{2^{-n}k\leavevmode\nobreak\ \colon\leavevmode\nobreak\ k\in\mathbb{Z}\}$ the one-dimensional grid of mesh size $2^{-n}$ for all $n\in\mathbb{N}$, we recall basic results of wavelet theory which we will use as starting point of our analysis, see [Mey93, Chapter 3] for the proof.
@@
As in [Ker23], we can replace the one-dimensional wavelets $\varphi,\hat{\varphi}$ with shifted wavelets $\varphi_{x},\hat{\varphi}_{x}$ for any $x\in\mathbb{R}$ without losing any properties of the wavelets. Thus we can assume that the wavelets $\phi,\hat{\phi}$ have a support in $[C,R]^{d}$ for some constants $0<C<R$ and that $a_{k}\neq 0$ only holds for $k\geq 0$.
@@
and it is easy to see that these equalities do hold for all $\psi\in C_{c}^{r}(\mathbb{R}^{d})$ as limits in $L^{2}(\mathbb{R}^{d})$. The limit $\mathbf{n}\to\infty$ stands for a limit along any sequence $(\mathbf{n}^{(k)})_{k\in\mathbb{N}^{d}}$ such that $n^{(k)}_{i}$ goes to $\infty$ for each $i\in[d]$. By [Ker23, Lemma 62], the above limits also hold in $C_{c}^{q}$ for $q<r$ as long as $r$ is big enough, as well as
@@
Note that the convergence speed is faster than one would expect: If one starts with any function $\phi$, one would expect the rates $2^{-\mathbf{n}/2}{\bm{\lambda}}^{-{\bm{1}}}$ simply from the localizations of $\phi^{\mathbf{n}}_{\mathbf{y}}$ and $\psi^{\bm{\lambda}}_{\mathbf{z}}$. However, the fact that $\phi^{\zeta}$ is orthogonal to polynomials of degree $\leq r$ in some $i\in\zeta$ allows us to get the additional factor $2^{-(r\mathbf{n})_{\zeta}}{\bm{\lambda}}^{-(r\mathbf{1})_{\zeta}}$. This is a more precise version of [Ker23, Lemma 61] (which itself is derived from [Hai14, Proposition 3.20]), where they only used the weaker condition, that all $\phi^{\zeta}$ for $\zeta\neq\emptyset$ eliminate multivariate polynomials of degree $\leq r$.

### v3

@@ Abstract
We extend the stochastic reconstruction theorem to a setting in which the underlying family of distributions satisfies natural conditions involving rectangular increments. This allows us to establish the well-posedness of a new class of mixed stochastic partial differential equations of hyperbolic type, which combine standard Walsh stochastic integration with Young products.
@@ 1.1 Stochastic reconstruction and hyperbolic SPDEs
where $\lambda\in(0,1]$, $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, $\psi_{\mathbf{x}}^{\lambda}(\cdot)=\lambda^{-d}\psi(\lambda^{-1}(\cdot-\mathbf{x}))$ is a rescaled and recentered version of a test function $\psi$ and $\lesssim$ denotes inequality up to a constant. Further analytic conditions were studied in [BL23, ZK23].
In this paper, we are interested in a generalization of this result where the underlying family $F$ is composed of random distributions depending on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ which are properly adapted to a multiparameter filtration =$\mathcal{F}=\{\mathcal{F}_{\mathbf{x}}\colon\mathbf{x}\in\mathbb{R}^{d}\}$ (see Definition 2.15). This setting is not novel in the literature. A previous result [Ker24] showed we can modify the conditions (1.1) into
@@
where $\delta>0$, $\|X\|_{m}=\mathbb{E}(|X|^{m})^{1/m}$ and $\mathbb{E}^{i}_{\mathbf{x}}(X)=\mathbb{E}[X|\mathcal{F}^{i}_{\mathbf{x}}]$ is the conditional expectation with respect to $\mathcal{F}^{i}_{\mathbf{x}}$, the filtration generated by the $i$-th component of $\mathbf{x}$. Remarkably, the uniqueness result holds even in the case $20>\gamma>-d/$ 2 provided that $\gamma+\delta>0$. Heuristically, we leverage the gain of regularity obtained by conditioning the germ by allowing a lower index of general regularity $\gamma$. This type of argument is not new in the rough analysis literature but it follows the same idea of the stochastic sewing lemma [Lê20], from which [Ker24] was inspired. Moreover, stochastic sewing arises as a special case of the stochastic reconstruction when $d=1$, see [Ker24, Section 5]. From a stochastic analysis perspective, stochastic reconstruction provides a new analytic framework to study Walsh integration [Wal86] with respect to a white noise $\xi$ over +$\mathbb{R}^{d}_{+}:=[0,+\infty)^{d}$ as a Young product, see [CZ20]. Recalling the existence of a measurable modification of $\xi$ in the space of distribution, for any adapted process $Y$ satisfying some weak regularity assumption one can check that the germ $(Y\cdot\xi)_{\mathbf{x}}(\psi):=Y_{\mathbf{x}}\xi(\psi)$ satisfies (1.2) and one obtains
@@
where the left-hand side is a standard Walsh integral, see [Ker24, Theorem 53].
@@
where ${\bm{\lambda}}=(\lambda_{1}\,,\ldots\,,\lambda_{d})\in(0,1]^{d}$, $\mathbf{x}=(x_{1}\,,\ldots,x_{d})$ $\mathbf{y}=(y_{1}\,,\ldots,y_{d})$ are in compact set, and $\psi_{\mathbf{y}}^{{\bm{\lambda}}}$ is a rescaled test function like in (2.7). Then there exist unique family of distributions $\mathcal{R}^{\theta}_{\mathbf{x}}(F)$ for each $\theta\subset\{1,\dots,d\}$ , such that $\mathcal{R}_{\mathbf{x}}^{\theta}(F)$ does not depend on $x_{i}$ for $i\in\theta$ and for any non empty subset $\eta\subset\theta\neq\emptyset$ one has the growth condition
@@
Unlike the statements in [CZ20, Ker24], in this statement we allow one to freeze certain variables $x_{i},i\notin\theta\subset\{1,\dots,d\}$ and only apply the reconstruction techniques for directions $i\in\theta$. As a result, we do not get a single distribution $\mathcal{R}(F)$, but a family of partial reconstructions $\mathcal{R}^{\theta}_{\mathbf{x}}(F)$. Let us illustrate these on example (1.3) in $d=2$. By the same arguments as [Ker24, Theorem 53],
@@
It should be noted that the proof of the stochastic reconstruction theorem is closer to the original proof of Martin Hairer than the later improvements [CZ20, BL23, ZK23]. This is because Martin Hairer’s proof is based on wavelet approximations, which approximate the target distribution $f$ with the sum
@@
where $\phi^{\mathbf{n}}_{\mathbf{y}}$ is a Daubechies-wavelet and $\Delta_{\mathbf{n}}$ is an increasingly fine mesh over $\mathbb{R}^{d}$, see Section 2. This particular structure allows to treat the convergence of these sums in the same way as [Lê20] by using a proper estimate on sums via a generalized BDG inequality, see Lemma 2.14. Moreover, wavelets also play an important role when we want to integrate the reconstruction to obtain a proper random field (see Proposition 4.7).
From the possibility of multiplying a Brownian sheet with a deterministic distribution via stochastic reconstruction in Theorem 1.1 we will be able to provide a full solution theory for a new type of mixed non-linear SPDEs
@@
where $\xi$ is a white noise over $\mathbb{R}_{+}^{d}$, $\sigma\colon\mathbb{R}\to\mathbb{R}$ is a generic non-linearity, $f\colon\Omega\times\mathbb{R}_{+}^{d}\to\mathbb{R}$ is a given random field, and $Z\colon\mathbb{R}_{+}^{d}\to\mathbb{R}$ is a deterministic Hölder field in $C^{\beta}$, see Section 3.1.
Equation (1.7) might seem unusual but when $d=2$ arises under different forms in new and old problems in PDE and SPDE theory: firstly, when $\sigma=0$ and $f\equiv 1$, the equation (1.7) becomes the signature kernel PDE, see [SCF+21], which plays an important role in the theory of signature Kernels, a possible theory to implement machine learning techniques with sequential data. In particular, equation (1.7) when $d=2$ represents a possible stochastic perturbation of the signature kernel PDE. Concerning the deterministic literature, equation (1.7) represents a stochastic perturbation of a special differential equation driven by a Hölder field on hyper-cubes, see [Har21], hence we decided to formulate the equation in a generic dimension $d$. Concerning the opposite case when $f=0$ the underlying SPDE is classical and it has been intensively studied, see e.g. [Cai72, NY89a, NY89b, MSS94] in dimension $d=2$. We also remark that the differential operator $\frac{\partial^{2}}{\partial x_{1}\partial x_{2}}u$ is naturally linked with the wave operator. Therefore equation (1.7) when $d=2$ could help us to understand path-wise well-posedness of stochastic wave equation in future investigations. From a mathematical perspective, equations such as (1.7) do not seem to have been addressed in the literature yet, and from a technical point of view, both classical deterministic techniques and techniques based on standard stochastic integration fail to establish well-posedness of this problem.
To give a rigorous meaning to equation (1.7), we recall the identification of stochastic integration with the reconstruction in (1.3) and we will describe the products in the right-hand side as reconstruction of germs of the simple form recalled before. Then we can write the equation into its integral form after applying an integration map
@@
Remark 1.5.
Recently, a new stochastic extension of the multiparameter sewing lemma was proposed in [BHK23] to study the effect of regularisation by noise phenomena on a similar SPDE. The techniques developed in that article are fundamental to the proof of the reconstruction theorem, which leads to Theorem 1.1 as a natural extension of stochastic multiparameter sewing in a distributional setting. Concerning the equation (1.7), using the $C^{\alpha,\delta}L_{m}$ spaces introduced in this paper, one can also solve it using the multiparameter stochastic sewing lemma, provided that the result of sewing lies in the appropriate spaces (see Remark 5.6), which we believe requires the same complexity as the results in Section 4. In contrast to the stochastic multiparameter sewing lemma, stochastic reconstruction has the advantage of not dealing with multiparameter delta operators $\delta^{\eta}$ (see Section 5), which become cumbersome in high dimensions. Moreover, it allows for a notion of general stochastic integration when there is no clear definition of a $2d$-parameter process to integrate (see Proposition 5.10). Finally, a formulation of this SPDE using a reconstruction theorem could be more helpful in extending techniques inspired by regularity structures to more general hyperbolic SPDEs. Further details on the relationship between our version of stochastic reconstruction and the stochastic multiparameter sewing lemma are studied in Section 5.
@@ 1.2 Outline of the paper
In Section 5, we show that the multiparameter stochastic sewing lemma from [BHK23] is a special case of our reconstruction theorem 3.12. This is shown in Proposition 5.4. It is based on the same construction, which connects sewing and reconstruction (see e.g. [BZ22]), as well as stochastic sewing and stochastic reconstruction (see [Ker24]), in dimension $d=1$.
@@ 1.3 Open extensions arising from our work
We conclude the introduction by outlining two natural extensions raised by the results we established. In our current formulation, we can only solve (1.8) for non-linearities of the form $f(\mathbf{x})u(\mathbf{x})$. One natural extension of the problem would be to consider a generic non-linearity $g(u)$ in (1.7). However, it seems that the structure of the spaces $C^{\alpha,\delta}L_{m}$ is not well suited to take generic compositions $g(u)$ for $u\in C^{\alpha,\delta}L_{m}$ into account. That is, even for smooth functions $g$, $g(u)$ might not belong to the same space as $u$. This stands in stark contrast to the deterministic case: In $d=2$, it is well-known that if $g\in C^{2}$ then $g(u)\in C^{\bm{\alpha}}$ for any $u\in C^{\bm{\alpha}}$ and ${\bm{\alpha}}=(\alpha_{1},\alpha_{2})\in(0,1)^{2}$ (see e.g. [QST07, Lemma 3.1]). Moreover, for any Hölder function $\sigma\in C^{\eta}$ for some $\eta\in(1,2)$, we still have that $\sigma(u)\in C^{\tilde{\alpha}}$ for a some $\tilde{\alpha}<\alpha$ (see [BHR24, Lemma 15]). Both of these results rely on the application of the classical Taylor formula to the rectangular increment
@@
In [FHL21], the authors developed a complete theory of rough stochastic differential equations of the form
$dY_{t}=b_{t}(Y_{t})+\sigma_{t}(Y_{t})dW_{t}+f_{t}(Y_{t})d\mathbf{X}_{t}$ (1.9)
[…]
@@ Acknowledgements
CB is supported by the ERC Starting Grant “LoRDeT.” He also acknowledges the support of TU Berlin, where he was employed when this project was initiated. HK is supported by the Research Unit FOR2402 at TU Berlin and the International Research Training Group 2544 of the Berlin Mathematical School. Both authors would like to thank the organizers of the conference “Structural Aspects of Signatures and Rough Paths” held at the Centre for Advanced Study in Oslo, where a preliminary version of these results was presented.
@@
$(\mathbf{y})^{\mathbf{x}}_{\theta}=\prod_{i\in\theta}y_{i}^{x_{i}}\,,$
@@ 2.2 Wavelet bases
To prove the reconstruction theorem in the rectangular increment setting we will have to use all the regularising effects of wavelet theory, going back to [Dau88]. First, we will recall some general facts about wavelet basis, mainly following [Hai14, Ker24, Mey95].
@@
where we define $\hat{\phi}^{\zeta,\mathbf{n}}_{\mathbf{x}}$ analogously. Using the notation $\Delta_{n}:=\{2^{-n}k\leavevmode\nobreak\ \colon\leavevmode\nobreak\ k\in\mathbb{Z}\}$ the one-dimensional grid of mesh size $2^{-n}$ for all $n\in\mathbb{N}$, we recall basic results of wavelet theory which we will use as starting point of our analysis, see [Mey95, Chapter 3] for the proof.
@@
As in [Ker24], we can replace the one-dimensional wavelets $\varphi,\hat{\varphi}$ with shifted wavelets $\varphi_{x},\hat{\varphi}_{x}$ for any $x\in\mathbb{R}$ without losing any properties of the wavelets. Thus we can assume that the wavelets $\phi,\hat{\phi}$ have a support in $[C,R]^{d}$ for some constants $0<C<R$ and that $a_{k}\neq 0$ only holds for $k\geq 0$.
@@
and it is easy to see that these equalities do hold for all $\psi\in C_{c}^{r}(\mathbb{R}^{d})$ as limits in $L^{2}(\mathbb{R}^{d})$. The limit $\mathbf{n}\to\infty$ stands for a limit along any sequence $(\mathbf{n}^{(k)})_{k\in\mathbb{N}^{d}}$ such that $n^{(k)}_{i}$ goes to $\infty$ for each $i\in[d]$. By [Ker24, Lemma 62], the above limits also hold in $C_{c}^{q}$ for $q<r$ as long as $r$ is big enough, as well as
@@
Note that the convergence speed is faster than one would expect: If one starts with any function $\phi$, one would expect the rates $2^{-\mathbf{n}/2}{\bm{\lambda}}^{-{\bm{1}}}$ simply from the localizations of $\phi^{\mathbf{n}}_{\mathbf{y}}$ and $\psi^{\bm{\lambda}}_{\mathbf{z}}$. However, the fact that $\phi^{\zeta}$ is orthogonal to polynomials of degree $\leq r$ in some $i\in\zeta$ allows us to get the additional factor $2^{-(r\mathbf{n})_{\zeta}}{\bm{\lambda}}^{-(r\mathbf{1})_{\zeta}}$. This is a more precise version of [Ker24, Lemma 61] (which itself is derived from [Hai14, Proposition 3.20]), where they only used the weaker condition, that all $\phi^{\zeta}$ for $\zeta\neq\emptyset$ eliminate multivariate polynomials of degree $\leq r$.